OpenStudy (anonymous):

The legs of a right triangle have lengths of 4 and 5 units. Find the length of the hypotenuse.

OpenStudy (anonymous):

Use Pythagorus' Theorem: \[c^2 = a^2+b^2\] Where c is the hypotenuse and a and b are the legs. You should end up with \[c^2 = 16+25\] Adding and taking the square root of both sides you get \[c = 6.40 \]

OpenStudy (amistre64):

i think grey is using british notation :)

OpenStudy (anonymous):

\[\sqrt{41}\]

OpenStudy (anonymous):

sorry >_<

OpenStudy (amistre64):

sqrt(41) looks good to me